haleyelizabeth2017:

CoCl2(aq)+2NaOH(aq)>Co(OH)2(s)+2NaCl(aq) How much 0.5 M NaOH solution in liters will completely precipitate the Co2+ in 1.6 L of 0.12 M CoCl2 solution?

haleyelizabeth2017:

@Tranquility sorry, me again. so far you're the only one who can explain this without 100% confusing me lol

Bob:

dang that's some problem

haleyelizabeth2017:

I've gotten the rest of stoichiometry down, just don't understand this XD

Yggdrasil:

You've calculated that you need 192ml(?) of NaOH(aq), yes? c = n/v 0.12 = n/1.6 n = 0.192 of CoCl2 we need 2x 0.192 moles of NaOH molecules = 0.384mol 0.5 = 0.348/V V = 0.192L or 192ml These calculations assume that the solutions are perfectly calibrated and not contaminated, also that all reactants will react.

Tranquility:

@haleyelizabeth2017 Let me know if you were able to understand @yggdrasil 's answer, otherwise I'd be more than happy to explain it further!

haleyelizabeth2017:

I was able to figure it out on my own after a while, but thank you!